# Differentially Private 2D Human Pose Estimation

### v2

@@ Abstract.
Human pose estimation (HPE) has become essential in numerous applications including healthcare, activity recognition, and human-computer interaction. However, the privacy implications of processing sensitive visual data present significant deployment barriers in critical domains. While traditional anonymization techniques offer limited protection and often compromise data utility for broader motion analysis, Differential Privacy (DP) provides formal privacy guarantees but typically degrades model performance when applied naively. In this work, we present the first differentially private 2D human pose estimation (2D-HPE) by applying Differentially Private Stochastic Gradient Descent (DP-SGD) to this task. To effectively balance privacy with performance, we adopt Projected DP-SGD (PDP-SGD), which projects the noisy gradients to a low-dimensional subspace. Additionally, we adapt TinyViT, a compact and efficient vision transformer for coordinate classification in HPE, providing a lightweight yet powerful backbone that enhances privacy-preserving deployment feasibility on resource-limited devices. Our approach is particularly valuable for multimedia interpretation tasks, enabling privacy-safe analysis and understanding of human motion across diverse visual media while preserving the semantic meaning required for downstream applications. Comprehensive experiments on the MPII Human Pose Dataset demonstrate significant performance enhancement with PDP-SGD achieving 78.48% PCKh@0.5 at a strict privacy budget ($\epsilon=0.2$), compared to 63.85% for standard DP-SGD. This work lays foundation for privacy-preserving human pose estimation in real-world, sensitive applications.
2D Human Pose Estimation, Differential Privacy
copyright: acmlicensedjournalyear: XXdoi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing Methodologies Activity recognition and understandingccs: Security and privacy Privacy protections
1. Introduction
@@
Human pose estimation (HPE) represents a fundamental multimedia interpretation task that transforms raw visual content into structured representations of human positioning and movements. This sophisticated conversion of visual data into anatomically meaningful keypoint configurations enable numerous high impact applications in healthcare, activity recognition, human-computer interaction, sports analysis and video games (Artacho and Savakis, 2020; Mao et al., 2022; Wang et al., 2022a; Lu et al., 2024; Li et al., 2022).
@@
As these multimedia systems are increasingly integrated into sensitive environments such as hospitals and homes, protecting user privacy has become a critical concern (202, [n. d.]). HPE systems that capture and process raw images pose significant privacy risks in two key ways: first, the raw images themselves contain sensitive personal information that could be exposed during collection and processing (Zheng et al., 2023); second, even trained models can inadvertently reveal information about individuals in training data when subjected to sophisticated privacy attacks such as model inversion or membership inference techniques (Geiping et al., 2020; Jegorova et al., 2023; Zakariyya et al., 2025). For instance, an adversary can exploit a model’s weights (Haim et al., 2022) or gradients (Hatamizadeh et al., 2023) to reconstruct distinctive physical characteristics of patients or sensitive contextual information, such as the patient’s home environment from the private training dataset (Jegorova et al., 2023). This reconstruction could potentially identify individuals with specific medical conditions and reveal that they received treatment at a particular facility during the model’s training period, thereby compromising both medical confidentiality and location privacy.
@@
Previous approaches to privacy preservation in HPE have focused primarily on data anonymization techniques, such as blurring, pixelation, and template-based shape modeling  (Ahmad et al., 2024; Ruiz-Zafra et al., 2023; Hesse et al., 2018). While these methods provide some level of privacy protection, they are often task-specific and can severely compromise the utility of the data for broader analysis. For instance, anonymization that removes facial features might preserve basic joint position information but destroy crucial clinical indicators needed for stress level assessment or abnormal motion pattern detection  (Barattin et al., 2023). Furthermore, these methods typically lack theoretical privacy guarantees and remain vulnerable to more sophisticated attacks, limiting their applicability in highly sensitive contexts  (Zakariyya et al., 2025). Moreover, these approaches do not address the inherent vulnerability of neural networks to memorization attacks that can reconstruct training data (Haim et al., 2022). This limitation significantly limits the potential to share models trained in sensitive multimedia datasets in the wider research community and clinical applications.
The inherent tension between improving model utility and ensuring robust privacy preservation represents a challenging research problem (Abbasi et al., 2025; Fang et al., 2023; Zhou et al., 2021) that has not been adequately explored in the context of HPE. Differential privacy (DP) offers a promising framework to address these concerns by providing mathematical guarantees against data leakage (Hardt and Talwar, 2010; Dwork, 2021; Abadi et al., 2016; Dwork et al., 2006). However, implementing DP in deep learning models, particularly through Differentially Private Stochastic Gradient Descent (DP-SGD)(Abadi et al., 2016), typically results in substantial performance degradation that can render models impractical for real-world applications  (Yu et al., 2019; De et al., 2022) . This degradation is problematic in vision tasks such as HPE, where the utility of the model relies on maintaining fine-grained spatial precision.
In this paper, we address the fundamental challenge of balancing the trade-off between privacy protection and model utility in HPE. Our work makes three significant contributions:
  • We present the first systematic application of DP-SGD to the Human Pose Estimation task. We empirically establish strong privacy-preserving baselines using DP-SGD on the MPII human pose dataset  (Rohrbach et al., 2015), and demonstrate the trade-off between utility and privacy in HPE models.
  • We successfully integrate Projected DP-SGD (PDP-SGD) to enhance the privacy-utility tradeoff in 2D-HPE. By projecting noisy gradients onto a low-dimensional subspace learned from a public dataset, we reduce noise amplification from DP and improve convergence, providing a more practical solution for tight privacy budgets.
  • We develop a novel lightweight and efficient HPE architecture based on TinyViT with a coordinate classification head, specifically designed to be compatible with differential privacy. This lightweight, DP-compatible architecture offers improved performance with reduced model size and serves as strong backbone for DP-optimized pose estimation, making privacy-preserving deployment more feasible on resource-limited devices.
Our experimental results on the MPII dataset  (Rohrbach et al., 2015) demonstrate that the proposed method achieves competitive performance compared to non-private baselines, with reasonable accuracy degradation at practical privacy budgets. Our projection-enhanced DP-SGD approach for HPE consistently improves pose estimation accuracy by up to 20% at strict privacy budgets ($\epsilon=0.2$), with the most substantial gains observed when using moderate gradient clipping thresholds (C=0.1, C=1.0). Fine-tuning pretrained models with carefully selected clipping norms further enhances performance over training from scratch. These results establish a practical pathway for deploying privacy-preserving HPE in sensitive domains. While our framework builds on existing techniques such as DP-SGD, we demonstrate its empirical viability for structured prediction tasks like HPE, addressing a gap in privacy-preserving computer vision.
2. Related Work
2.1. Markerless 2D Human Pose Estimation
@@
Markerless 2D-HPE identifies anatomical keypoints in images without physical markers, thus playing a fundamental role in human motion analysis for several applications in healthcare and activity recognition  (Deligianni et al., 2019; Bondugula et al., 2023). Although traditional heatmap methods originally based on CNN architectures (Tompson et al., 2014; Artacho and Savakis, 2020; Debnath et al., 2018; Zhang et al., 2020; Kamel et al., 2020; Wang et al., 2022b; Xiao et al., 2018) and later leveraging vision-based transformers (Li et al., 2021b, d; Ma et al., 2022; Yuan et al., 2021; Xu et al., 2022) achieved state-of-the-art performance, they suffer from quantization errors and usually result in large cumbersome networks. Regression approaches offer faster, end-to-end solutions but with reduced accuracy (Toshev and Szegedy, 2014; Zhou et al., 2019; Nie et al., 2019; Li et al., 2021a). Recently, a coordinate classification approach has been proposed to address these limitations by treating pose estimation as a classification problem with discretized coordinates (Li et al., 2022). This method achieved state-of-the-art performance while maintaining computational efficiency. Knowledge distillation techniques (Ye et al., 2023; Li et al., 2021c; Zhang et al., 2023; Bhargav Sivangi, 2024) have been also applied to enhance the efficiency of HPE with significantly less model parameters and inference speed by transfering knowledge from large heatmap based models to smaller architectures.
2.2. Privacy Preservation in 2D-Human Pose Estimation
@@
Protecting the privacy of users is of paramount importance in all real-world applications and a challenging task for human pose estimation, which depends on high-quality images. Both trained models and training datasets can become vulnerable targets for attacks in modern multimedia systems. Recent research has demonstrated that adversaries can successfully reconstruct substantial portions of private training data solely by analyzing the parameters of a trained neural network (Haim et al., 2022; Jegorova et al., 2023). For these reasons, in critical applications, data sharing initiatives provide limited information, such as the location of limb joints and depth images (Janet Pigueiras-del Real, 2024). This restricted access prevents the analysis of crucial clinical indicators that require body shape or facial information to detect levels of stress and identify abnormal motion patterns. In practice, most platforms implement rudimentary privacy protection based on face de-identification with blurring and pixelation (Ruiz-Zafra et al., 2023). Some approaches to protecting people’s privacy have evolved from basic techniques to more sophisticated methods, although significant challenges persist. Early work by (Hesse et al., 2018) employed skin removal and template-based shape modeling, which compromised clinical utility despite privacy benefits. Recently, more sophisticated method have been developed. In (Hinojosa et al., 2021), an end-to-end framework is proposed that integrates an optimized optical encoder with a CNN decoder, incorporating a visual privacy protection layer to degrade private attributes while preserving essential features. These methods lack theoretical privacy guarantees and remain vulnerable to reconstruction attacks (Zakariyya et al., 2025). Furthermore, they are fine-tuned for a specific task, which implies that extracting additional useful information from the data might not be possible.
2.3. Utility vs privacy with DP optimization
DP-SGD is one of the most common methods in developing privacy-preserved deep learning models because of the strong privacy guarantees compared to data-independent methods and its ability to scale to large datasets (Hardt and Talwar, 2010; Dwork, 2021). In DP-SGD the process involves clipping the gradients first to bound the maximum norm of individual gradients and subsequently adding Gaussian noise. Clipping the gradients ensures that the gradients’ sensitivity is bounded so that any single data point would have very limited influence (Lebensold et al., 2024). In this way, DP is maintained by preventing a single data point disproportionally influencing the deep learning model.
However, there is a trade-off between privacy and utility in DP settings that can compromise the ability to develop privacy-preserved HPE algorithms that would be useful in practice (Abbasi et al., 2025). The performance of DP-SGD has been shown to depend on the smoothness of the loss function (Wang et al., 2020) and the dimension of the gradient (Fang et al., 2023; Zhou et al., 2021). It has also been shown that if the loss function lacks Lipschitz continuity, the performance of DP-SGD critically depends on carefully selecting an appropriate clipping threshold; otherwise, performance will not improve significantly, regardless of the amount of training data or iterations (Fang et al., 2023). To our knowledge, in the context of preserving the privacy of the models in HPE, there is no work that studies the utility-privacy trade-off.
3. Methodology
Towards our goal of exploring the intersection of privacy and utility in HPE, we design a modular framework for differentially private 2D HPE. Our approach assumes access to a public dataset, used for pretraining, and a private dataset on which the model is subjected to formal DP constraints. At the core, we employ DP-SGD with calibrated clipping and noise addition, aligning with the structured nature of pose outputs (Sec. 3.1). To address the degradation of utility under DP-SGD, we incorporate Projected DP-SGD (PDP-SGD) (Zhou et al., 2021), which projects the noisy gradients associated with private set onto a gradient subspace derived from a subset of public data (Sec. 3.3). Finally, we instantiate our framework using a novel DP-compatible architecture based on the compact TinyViT backbone with a coordinate classification head (Sec. 3.4).
3.1. Differential Privacy (DP)
@@
To ensure privacy preservation in any application, it is essential to establish a quantifiable definition of privacy. DP provides a rigorous mathematical framework to quantify the privacy guarantees of a system  (Hardt and Talwar, 2010). Specifically, in the context of databases, DP measures how much the inclusion or exclusion of a single data point can influence the output of a randomized algorithm  (Dwork, 2021).
@@
Consider two neighboring datasets, i.e., datasets that differ by only a single sample. The level of DP achieved by a randomized algorithm $\mathcal{M}$ is defined as follows:
@@
Definition 1: $(\epsilon,\delta)$-DP A randomized algorithm $\mathcal{M}$ with domain $\mathcal{D}$ and range $\mathcal{R}$ is said to be $(\epsilon,\delta)$-DP if, for any subset $S\subseteq\mathcal{R}$ and for any neighboring datasets $d,d^{\prime}\in\mathcal{D}$, the following condition holds:
(1) $\mathbb{P}[\mathcal{M}(d)\in S]\leq e^{\epsilon}\cdot\mathbb{P}[\mathcal{M}(d^{\prime})\in S]+\delta$
@@
This definition ensures that the probability of obtaining a particular output does not change significantly when a single data point is modified, thereby providing a formal measure of privacy protection  (Dwork, 2021). In HPE, this definition implies that the inclusion or exclusion of a single training image, potentially representing an identifiable subject, should not significantly affect the model’s predictions across downstream tasks. DP trained model ensures no particular subject’s identity or appearance can be inferred from model’s parameters and predictions mitigating risks such as membership inference(Shokri et al., 2017) or reconstruction attacks(Zhu et al., 2019).
Sensitivity and the Gaussian Mechanism: The most common method to enforce DP in an algorithm is by adding noise sampled from a Gaussian distribution. The amount of noise required is typically controlled by a parameter known as sensitivity, denoted as $\Delta_{G}$, which quantifies how much the output of a function can change when a single data point is modified  (Lebensold et al., 2024).
Definition 2: L1 Sensitivity For any function $G:\mathbb{R}\rightarrow\mathbb{R}^{*}$, the L1 sensitivity is the smallest value $\Delta_{G}$ such that for all neighboring datasets $d,d^{\prime}$:
(2) $\left\|G(d)-G(d^{\prime})\right\|_{L1}\leq\Delta_{G}$
Intuitively, sensitivity measures the maximum influence that a single data point can have on the query output. A higher sensitivity value requires adding more noise to ensure privacy. In the Gaussian mechanism, the inputs consist of the randomized algorithm $\mathcal{M}$, its sensitivity $\Delta_{M}$, and the noise scale parameter $\sigma$. The perturbation applied to the output follows a Gaussian distribution:
(3) $Gaussian\_Mechanism(\mathcal{M},\Delta_{M},\sigma)\stackrel{{\scriptstyle\Delta}}{{=}}\mathcal{M}+\mathcal{N}(0,\Delta_{M}^{2}\cdot\sigma^{2})$
where $\mathcal{N}(0,\Delta_{M}^{2}\cdot\sigma^{2})$ represents a Gaussian distribution with mean 00 and variance $\Delta_{M}^{2}\cdot\sigma^{2}$. The level of privacy protection is directly influenced by the noise scale, which is determined by $\sigma$.
Based on the analysis in (Abadi et al., 2016), for any $(\epsilon,\delta)\in(0,1)$, the Gaussian mechanism achieves $(\epsilon,\delta)$-differential privacy when:
(4) $\sigma=\frac{\sqrt{2\ln(1.25/\delta)}}{\epsilon}$
This formulation provides a practical approach to achieving differential privacy guarantees while controlling the trade-off between privacy and utility.
3.2. Differentially Private Stochastic Gradient Descent (DP-SGD)
@@
A common approach to ensure DP during model is training is DP-SGD, which enforces an $(\epsilon,\delta)$-DP guarantee on gradient updates (Abadi et al., 2016; Dupuy et al., 2022; Boenisch et al., 2024). This mechanism involves clipping gradients to a fixed threshold ($C$) and adding Gaussian noise calibrated based on desired privacy budget ($\epsilon$, $\delta$). This process limits the influence of any single training example on the model update, thereby providing formal privacy protection (Kong and Munoz Medina, 2023). The model is trained iteratively using mini-batches of input data with the SGD algorithm; this procedure is termed DP-SGD. As described in Algorithm 1, it requires an input dataset $\mathcal{D}$, clipping norm $C$ and noise scale $\sigma$, which is determined by the privacy guarantee parameters $(\epsilon,\delta)$. Lower values of $\epsilon$ and $\delta$ indicates stronger privacy protection and better resistance to privacy attacks. The DP-SGD procedure in Algorithm 1 follows the standard mini-batch training approach, with the addition of lines 6, and 9, which implement gradient clipping, aggregation of the clipped gradients, and noise addition, respectively.
In our setting, we apply DP-SGD to train a TinyViT based 2D HPE model on MPII dataset. Due to the inherent structured nature of keypoint predicition task, a straight forward application of DP-SGD leads to severe degradation in performance. We empirically tune the clipping $C$ at different rates to identify configurations that offer an effective privacy-utility tradeoff. Additionally, we employ finetuning strategy in this setting, which allows us to recover utility under strict privacy budgets(Yu et al., 2021).
Algorithm 1 DP-SGD Procedure
Input: Dataset $\mathcal{D}=\{x_{1},\dots,x_{n}\}$, loss function $\mathcal{L}$, learning rate $\eta$, clipping norm $C$, noise scale $\sigma$, batch size $B$, number of iterations $T$
Output: Model parameters $\theta_{T}$
1:  Initialize model parameters $\theta_{0}$
2:  for $t=1$ to $T$ do
3:     Sample a random mini-batch $\mathcal{B}_{t}\subset\mathcal{D}$ of size $B$
4:     for each $x_{i}\in\mathcal{B}_{t}$ do
5:        Compute the gradient $\mathbf{g}_{i}=\nabla_{\theta}\mathcal{L}(\theta_{t-1},x_{i})$
6:        Clip the gradient: $\mathbf{\tilde{g}}_{i}=\mathbf{g}_{i}/\max\left(1,\frac{\|\mathbf{g}_{i}\|_{2}}{C}\right)$
7:     end for
8:     Aggregate the clipped gradients: $\mathbf{\bar{g}}=\frac{1}{B}\sum_{i\in\mathcal{B}_{t}}\mathbf{\tilde{g}}_{i}$
9:     Add noise: $\mathbf{\hat{g}}=\mathbf{\bar{g}}+\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I})$
10:     Update the model: $\theta_{t}=\theta_{t-1}-\eta\mathbf{\hat{g}}$
11:  end for
12:  return  $\theta_{T}$
3.3. Gradient Subspace Identification
@@
Our objective in this section is to improve the privacy-utility trade-off in HPE by reducing the impact of noise injection while maintaining the same privacy guarantees. The key insight driving our approach is that gradient updates during training tend to concentrate within a low-dimensional subspace rather than utilizing the full parameter space. By identifying and projecting noisy gradients onto this informative subspace, we filter out less relevant directions. In this way, we preserve the signal quality of gradient updates while adhering to DP constraints (Zhou et al., 2020), and removing privacy-sensitive information that is not strictly required to minimize the loss, such as facial features. To estimate the intrinsic structure of the gradient space, we employ a small auxiliary public dataset $S_{pub}$, which is drawn from a similar distribution as that of private training set. This subset is used to estimate the principal subspace of the gradient covariance. Given the model parameters $w\in\mathbb{R}^{p}$, the second moment matrix of gradients over $S_{pub}$ is calculated as:
(5) $M(w)=\frac{1}{m}\sum_{i=1}^{m}\nabla l(w,\tilde{z_{i}})\nabla l(w,\tilde{z_{i}})^{T}$

### v3

@@ Abstract
Human pose estimation (HPE) has become essential in numerous applications including healthcare, activity recognition, and human-computer interaction. However, the privacy implications of processing sensitive visual data present significant deployment barriers in critical domains. While traditional anonymization techniques offer limited protection and often compromise data utility for broader motion analysis, Differential Privacy (DP) provides formal privacy guarantees but typically degrades model performance when applied naively. In this work, we present the first comprehensive framework for differentially private 2D human pose estimation (2D-HPE) by applying Differentially Private Stochastic Gradient Descent (DP-SGD) to this task. To effectively balance privacy with performance, we adopt Projected DP-SGD (PDP-SGD), which projects the noisy gradients to a low-dimensional subspace. Next, we incorporate Feature Differential Privacy(FDP) to selectively privatize only sensitive features while retaining public visual cues. Finally, we propose a hybrid feature-projective DP framework that combines both approaches to balance privacy and accuracy for HPE. We evaluate our approach on the MPII dataset across varying privacy budgets, training strategies, and clipping norms. Our combined feature-projective method consistently outperforms vanilla DP-SGD and individual baselines, achieving up to 82.61% mean PCKh@0.5 at $\epsilon=0.8$, substantially closing the gap to the non-private performance. This work lays foundation for privacy-preserving human pose estimation in real-world, sensitive applications.
Introduction
@@
Human pose estimation (HPE) represents a fundamental task that transforms raw visual content consisting of humans into structured representations of human positioning and movements. This sophisticated conversion of visual data into anatomically meaningful keypoint configurations enables numerous high impact applications in healthcare, activity recognition, human-computer interaction, sports analysis, and video games (Artacho and Savakis 2020; Mao et al. 2022; Wang et al. 2022b; Lu et al. 2024; Li et al. 2022).
@@
As these human motion analysis systems are increasingly deployed into sensitive environments such as hospitals and homes, protecting user privacy has become a critical concern (Martinez-Martin et al. 2021). HPE systems that capture and process raw images pose significant privacy risks in two key ways: first, the raw images themselves contain sensitive personal information that could be exposed during collection and processing (Zheng et al. 2023); second, even trained models can inadvertently reveal information about individuals in training data when subjected to sophisticated privacy attacks such as model inversion or membership inference techniques (Geiping et al. 2020; Jegorova et al. 2023; Zakariyya et al. 2025). For instance, an adversary can exploit a model’s weights (Haim et al. 2022) or gradients (Hatamizadeh et al. 2023) to reconstruct distinctive physical characteristics of patients or sensitive contextual information, such as the patient’s home environment from the private training dataset (Jegorova et al. 2023). This reconstruction could potentially identify individuals with specific medical conditions and reveal that they received treatment at a particular facility during the model’s training period, thereby compromising both medical confidentiality and location privacy.
@@
Previous approaches to privacy preservation in HPE have focused primarily on data anonymization techniques, such as blurring, pixelation, and template-based shape modeling  (Ahmad, Morerio, and Del Bue 2024; Ruiz-Zafra et al. 2023; Hesse et al. 2018). While these methods provide some level of privacy protection, they are often task-specific and can severely compromise the utility of the data for broader analysis. For instance, anonymization that removes facial features might preserve basic joint position information but destroy crucial clinical indicators needed for stress level assessment or abnormal motion pattern detection  (Barattin et al. 2023). Furthermore, these methods typically lack theoretical privacy guarantees and remain vulnerable to more sophisticated attacks, limiting their applicability in highly sensitive contexts  (Zakariyya et al. 2025). Moreover, these approaches do not address the inherent vulnerability of neural networks to memorization attacks that can reconstruct training data (Haim et al. 2022). This limitation significantly limits the potential to share models trained in sensitive multimedia datasets in the wider research community and clinical applications. The inherent tension between improving model utility and ensuring robust privacy preservation represents a challenging research problem (Abbasi, Mori, and Saracino 2025; Fang et al. 2023; Zhou, Wu, and Banerjee 2021) that has not been adequately explored in the context of HPE. Differential privacy (DP) offers a promising framework to address these concerns by providing mathematical guarantees against data leakage (Hardt and Talwar 2010; Dwork 2021; Abadi et al. 2016; Dwork et al. 2006). However, implementing DP in deep learning models, particularly through Differentially Private Stochastic Gradient Descent (DP-SGD)(Abadi et al. 2016), typically results in substantial performance degradation that can render models impractical for real-world applications  (Yu et al. 2019; De et al. 2022) . This degradation is problematic in vision tasks such as HPE, where the utility of the model relies on maintaining fine-grained spatial precision.
In this work, we present the first comprehensive study of differential privacy for the 2D-HPE task. We demonstrate that directly applying DP-SGD to 2D-HPE models leads to significant degradation in utility due to the fine-grained nature of keypoint prediction. To address the privacy-utility trade-off in 2D-HPE, we explore multiple complementary strategies. First, we adopt a projection based DP-SGD approach that leverages the low-dimensional structure of the gradient space. By projecting the noisy gradients onto a principal subspace from a small public set, we mitigate privacy-induced noise without compromising the DP guarantees. Second, we incorporate Feature Differential Privacy (FDP), a privacy framework that relaxes differential privacy on public features, while rigorously protects sensitive features, thereby reducing the amount of calibrated noise required. Finally, we propose a hybrid strategy that integrates both projection and FDP effectively, combining their strengths.
To summarise, our core contributions are as follows:
  • We present the first systematic exploration of differentially private learning for 2D human pose estimation, establishing baseline performance under various privacy regimes.
  • We propose a projection-based DP-SGD mechanism adopted for pose estimation, which improves performance by restricting gradient updates to low-dimensional subspaces that capture the most informative directions.
  • We adapt and extend the Feature Differential Privacy (FDP) framework to the HPE domain, allowing selective private updates only for sensitive features. This approach automatically protects both human subjects and their surrounding spatial context without requiring manual feature curation. Additionally, we integrate this with the projection method to yield a unified feature-projective DP approach demonstrating consistent best privacy-utility trade-off across various settings.
  • We perform a thorough evaluation across multiple DP regimes, training strategies, and gradient clipping thresholds on the MPII dataset, demonstrating the effectiveness of the proposed method in improving the privacy-utility trade-off.
Related Work
2D Human Pose Estimation and Privacy
@@
Markerless 2D-HPE identifies anatomical keypoints in images without physical markers, thus playing a fundamental role in human motion analysis for several applications in healthcare and activity recognition  (Deligianni, Guo, and Yang 2019; Bondugula, Udgata, and Sivangi 2023). Although traditional heatmap methods originally based on CNN architectures (Artacho and Savakis 2020; Kamel et al. 2020; Wang et al. 2022a; Xiao, Wu, and Wei 2018) and later leveraging vision-based transformers (Li et al. 2021b, c; Xu et al. 2022) achieved state-of-the-art performance, they suffer from quantization errors and usually result in large cumbersome networks. Regression approaches offer faster, end-to-end solutions but with reduced accuracy (Toshev and Szegedy 2014; Nie et al. 2019; Li et al. 2021a). Coordinate classification approach addresses these limitations by treating pose estimation as a classification problem with discretized coordinates (Li et al. 2022). This method achieved state-of-the-art performance while maintaining computational efficiency. Knowledge distillation techniques (Ye et al. 2023; Li et al. 2021d; Zhang et al. 2023; Bhargav Sivangi 2024) have been also applied to enhance the efficiency of HPE with significantly less model parameters and inference speed by transfering knowledge from large heatmap based models to smaller architectures.
@@
Protecting the privacy of users is of paramount importance in all real-world applications and a challenging task for human pose estimation, which depends on high-quality images. Recent research works(Haim et al. 2022; Zhu, Liu, and Han 2019) have demonstrated that adversaries can successfully reconstruct substantial portions of private training data solely by analyzing the parameters and gradients of a trained neural network. For these reasons, in critical applications, data sharing initiatives provide limited information, which prevents the analysis of crucial clinical indicators that require body shape or facial information to detect levels of stress and identify abnormal motion patterns. In practice, most platforms implement rudimentary privacy protection based on face de-identification with blurring and pixelation (Ruiz-Zafra et al. 2023). Some approaches to protecting people’s privacy have evolved from basic techniques to more sophisticated methods, although significant challenges persist. Early work by (Hesse et al. 2018) employed skin removal and template-based shape modeling, which compromised clinical utility despite privacy benefits. (Hinojosa, Niebles, and Arguello 2021) proposed an end-to-end framework that integrates an optimized optical encoder with a CNN decoder, incorporating a visual privacy protection layer to degrade private attributes while preserving essential features. These methods lack theoretical privacy guarantees and remain vulnerable to adversarial attacks (Zakariyya et al. 2025). Furthermore,traditional methods like blurring or masking significantly degrade data quality, reducing their effectiveness in downstream tasks like instance segmentation and pose estimation. Recent image anonymization techniques (Hukkelås and Lindseth 2023a; Hukkelås et al. 2023) have been developed to address privacy concerns, employing realistic GAN-based approaches such as to generate anonymized human figures and faces without severe visual distortion. Although realistic anonymization improves data utility compared to traditional obfuscation, it still introduces artifacts or unnatural appearances due to challenges like limited pose detection accuracy and contextual mismatches. Consequently, even advanced realistic anonymization techniques cannot yet fully replace real, non-anonymized data for training robust computer vision models (Hukkelås and Lindseth 2023b).
Utility vs Privacy with DP optimization
DP-SGD is one of the most common methods in developing privacy-preserved deep learning models because of the strong privacy guarantees compared to data-independent methods and its ability to scale to large datasets (Hardt and Talwar 2010; Dwork 2021). In DP-SGD the process involves clipping the gradients first to bound the maximum norm of individual gradients and subsequently adding Gaussian noise. Clipping the gradients ensures that the gradients’ sensitivity is bounded so that any single data point would have very limited influence (Lebensold, Precup, and Balle 2024). In this way, DP is maintained by preventing a single data point disproportionally influencing the deep learning model. However, there is a trade-off between privacy and utility in DP settings that can compromise the ability to develop privacy-preserved HPE algorithms that would be useful in practice (Abbasi, Mori, and Saracino 2025). The performance of DP-SGD has been shown to depend on the smoothness of the loss function (Wang et al. 2020) and the dimension of the gradient (Fang et al. 2023; Zhou, Wu, and Banerjee 2021). It has also been shown that if the loss function lacks Lipschitz continuity, the performance of DP-SGD critically depends on carefully selecting an appropriate clipping threshold; otherwise, performance will not improve significantly, regardless of the amount of training data or iterations (Fang et al. 2023). Recent research has sought to improve the utility-privacy trade-off by relaxing traditional differential privacy. (Shi et al. 2021) enhanced the utility of DP by introducing Selective Differential Privacy, which protects only sensitive tokens within language models, while leaving non-sensitive elements unperturbed. (Mahloujifar et al. 2025) proposed Feature Differential Privacy, a generalized DP framework that explicitly categorizes features as either protected or public, enabling targeted noise application that enhances utility. Both methods demonstrate that targeted protection of sensitive attributes substantially mitigates the privacy-utility trade-off compared to classical DP.
Figure 1: Overview of our private HPE pipeline coupling a TinyViT based backbone with Coordinate Classification Keypoint head. The Public feature batch $B_{pub}$ is generated from the private batch $B_{priv}$ using $\psi$ both of which are given as input in a single iteration. Additionally, a public image set $B_{pub}^{proj}$ independent of $B_{priv}$ is used to calculate public gradients for projection at specified intervals. Red Arrow indicates the propogation of private gradients and Green Arrow indicates propagation of public feature gradients. Blue Dotted Arrow indicates the propogation of cumulative denoised gradient from Feature Projective DP.
Methodology
Preliminaries: Differentially Private Stochastic Gradient Descent (DP-SGD)
@@
Privacy-preserving machine learning requires a rigorous mathematical framework to quantify privacy guarantees. DP provides such a framework by measuring in the context of databases how much the inclusion or exclusion of a single data point can influence the output of a randomized algorithm  (Dwork 2021; Hardt and Talwar 2010). This comparison allows us to bound the information leakage about any individual data point.
@@
Consider two neighboring datasets-identical except for a single sample. The level of DP ensured by a randomized algorithm $\mathcal{M}$ is provided by the following definition.
@@
Definition 1: $(\epsilon,\delta)$-Differential Privacy A randomized algorithm $\mathcal{M}$ with domain $\mathcal{D}$ and range $\mathcal{R}$ is said to be $(\epsilon,\delta)$-DP if, for any subset $S\subseteq\mathcal{R}$ and for any neighboring datasets $d,d^{\prime}\in\mathcal{D}$, the following condition holds:
$\mathbb{P}[\mathcal{M}(d)\in S]\leq e^{\epsilon}\cdot\mathbb{P}[\mathcal{M}(d^{\prime})\in S]+\delta$ (1)
In this definition, $\epsilon$ (privacy budget) controls the strength of the privacy guarantee. Smaller values of $\epsilon$ provide stronger privacy. The parameter $\delta$ represents the probability that the privacy guarantee fails and is typically set to be cryptographically small.
@@
In the context of HPE, differential privacy ensures that the inclusion or exclusion of a single training image, which potentially contains identifiable biometric information, does not significantly affect the model’s learned parameters or predictions. Therefore, DP trained models provide a formal measure of privacy protection (Dwork 2021), effectively mitigating risks from various attacks, including membership inference(Shokri et al. 2017) or reconstruction attacks(Zhu, Liu, and Han 2019).
DP-SGD
@@
A common approach for ensuring differential privacy during neural network training is DP-SGD, which enforces an $(\epsilon,\delta)$-DP guarantee on gradient updates (Abadi et al. 2016; Dupuy et al. 2022; Boenisch et al. 2024). This mechanism involves clipping gradients to a fixed $L2$-norm threshold ($C$) and adding Gaussian noise calibrated based on desired privacy budget ($\epsilon$, $\delta$). This process ensures that no single training sample can disproportionately influence the model update (Kong and Munoz Medina 2023).
Lightweight Architecture for 2D-HPE
For our 2D-HPE models, we adopt TinyViT (Wu et al. 2022) as the backbone. It is a small sized four-stage efficient hierarchical vision transformer which is well suited for resource-constrained vision tasks. The model adopts a multi-stage architecture where in the spatial resolution is progressively reduced and the feature representation expands. TinyViT follows a hybrid architectural design containing convolutional layers at the initial stages followed by self-attention mechanisms. Unlike standard ViT models, TinyViT employs a two-layer convolutional embedding. In the first stage of the network, it employs MBConv (Howard et al. 2019) blocks from MobileNetV2 to efficiently learn the low-level representation. The last three stages consists of transformer blocks hierarchically. Each stage consists of multi-head-self-attention (MHSA) layers, feed forward network (FFN) and $3$ x $3$ depthwise convolutions between the MHSA and FFN layers.
For keypoint localization, we augment the TinyViT backbone model with a coordinate classification output stage (Li et al. 2022). Given an input image $I\in\mathbb{R}^{C\times H\times W}$ and a ground truth keypoint $p_{i}=(x_{i},y_{i})$ for the $i^{\text{th}}$ joint, the continuous coordinates are quantized into discrete bins via a splitting factor $k\geq 1$. Formally, the quantized coordinates are computed as:
where $\lfloor\cdot\rfloor$ denotes the rounding operation. This binning reduces quantization error while preserving high localization precision. The complete architecture is depicted in Figure 1.
Projection Based DP-SGD
@@
Training dynamics in deep networks exhibit intrinsic low-dimensional structure, where meaningful gradient updates concentrate within a subspace significantly smaller than the full parameter space. We leverage this by identifying and projecting noisy gradients onto informative subspaces, filtering out less relevant directions while preserving signal quality under differential privacy constraints. In this way, we preserve the signal quality of gradient updates while adhering to DP constraints (Zhou, Wu, and Banerjee 2020). To estimate the intrinsic structure of the gradient space, we employ a small auxiliary public dataset $S_{pub}$, which is drawn from a similar distribution as that of private training set. This subset is used to estimate the principal subspace of the gradient covariance. Given the model parameters $w\in\mathbb{R}^{p}$, the second moment matrix of gradients over $S_{pub}$ is calculated as:
$M(w)=\frac{1}{m}\sum_{i=1}^{m}\nabla l(w,\tilde{z_{i}})\nabla l(w,\tilde{z_{i}})^{T}$ (2)